OpenStudy (taylev105):

If (4^4)x = 4^32, what is the value of x?

OpenStudy (taylev105):

@iambatman ,@iGreen ,@ganeshie8 help fan and medal

OpenStudy (taylev105):

@iGreen @ganeshie8

OpenStudy (taylev105):

@johnweldon1993 @Jhannybean

OpenStudy (taylev105):

omg anyone I need help????

OpenStudy (johnweldon1993):

Well \[\large (4^4)^x = 4^{32}\] we know that when we have exponents seperated by parenthesis...we multiply... so we need something times 4..equals 32

OpenStudy (taylev105):

8

OpenStudy (johnweldon1993):

Right...x would equal 8 \[\large (4^4)^8 = 4^{4\times 8} = 4^{32}\]

OpenStudy (anonymous):

x=8

OpenStudy (taylev105):

now u come @iGreen

OpenStudy (igreen):

Sorry, I was helping someone else..lol.

OpenStudy (taylev105):

its ok lol

OpenStudy (igreen):

Well yeah, all you had to do was divide. \((4^4)^x = 4^{32} \rightarrow 4^{4x=32}\) Divide 4 to both sides: \((4^4)^x = 4^{32} \rightarrow 4^{4x=32} \rightarrow 4^{x = 8} \rightarrow 4^8\)

OpenStudy (taylev105):

yea .....